# PLaID++: A Preference Aligned Language Model for Targeted Inorganic Materials Design

### v1

@@ 6 Discussion
Limitations & Future Work   While PLaID++ achieves state-of-the-art stability and S.U.N. rates, it has several limitations that point toward promising avenues for future work. We have fine-tuned and evaluated exclusively on the MP-20 dataset ($\sim$45K structures), and scaling to larger, more diverse datasets such as Alexandria ($\sim$2.6M structures) (Schmidt et al., 2024) or LeMaterial ($\sim$6.7M structures) (Siron et al., 2024) may further broaden the model’s coverage of chemical space. Our experiments also focus on unconditional and space-group-conditioned generation, but extending generation to other objectives such as band-gap optimization or ionic conductivity targets would demonstrate wider practical utility. We also only explored Direct Preference Optimization (DPO) as our RL method; investigating alternative algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (Shao et al., 2024) might uncover different stability–diversity trade-offs.
Impact Statement   Our approach has the potential to dramatically speed up the design of novel materials for renewable energy, electronics, and carbon-capture applications—paving the way for more efficient, sustainable technologies that benefit society. At the same time, deploying generative models carries risks, including the synthesis of dangerous compounds and uneven access to these tools.

### v2

@@ 6 Discussion
Limitations & Future Work   While PLaID++ achieves state-of-the-art stability and S.U.N. rates, it has several limitations that point toward promising avenues for future work. We have fine-tuned and evaluated exclusively on the MP-20 dataset ($\sim$45K structures), and scaling to larger, more diverse datasets such as Alexandria ($\sim$2.6M structures) (Schmidt et al., 2024) or LeMaterial ($\sim$6.7M structures) (Siron et al., 2024) may further broaden the model’s coverage of chemical space. Our experiments also focus on unconditional and space-group-conditioned generation, but extending generation to other objectives such as band-gap optimization would demonstrate wider practical utility. We also only explored Direct Preference Optimization (DPO) as our RL method; investigating alternative algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (Shao et al., 2024) might uncover different stability–diversity trade-offs.
Impact Statement   Our approach has the potential to accelerate the discovery of novel materials for renewable energy, electronics, and carbon-capture applications—paving the way for technologies that benefit society. However, deploying generative models carries risks, including the synthesis of dangerous compounds and inequitable access to these tools.